OpenStudy (anonymous):

When solving for a in -8+square route of (5a-5)=-3, do I square -8 when i get rid of the square route, and if so does it become 64 or -64?

OpenStudy (anonymous):

add 8 first!

OpenStudy (anonymous):

otherwise you will have a mess. \[-8+\sqrt{3a-5}=-3\] \[\sqrt{3a-5}=5\] etc

OpenStudy (anonymous):

first add 8, then square, then solve

OpenStudy (anonymous):

oh thanks i completely forgot about adding the 8!

OpenStudy (anonymous):

then check

OpenStudy (anonymous):

oh damn ignore my answer. i used \[3a-5\] instead of \[5a-5\] sorry

OpenStudy (anonymous):

that's fine i can solve it now that i know to add the 8. thanks!

OpenStudy (anonymous):

\[\sqrt{5a-5}=5\] \[5a-5=25\] \[5a=30\] \[a=6\] is more like it